OpenStudy (anonymous):

Can someone help me simplify (sec^2x-tan^2x)/(cscx)? I don't understand these problems

OpenStudy (imstuck):

Actually, JungHyunRan, it doesn't look like the csc x is squared in the denominator, so that would change things a bit. It would be equal to sinx instead of sin^2 x.

OpenStudy (anonymous):

oops.. I got a big error

OpenStudy (anonymous):

\(\frac { \sec ^{ 2 }{ x } -\tan ^{ 2 }{ x } }{ \sec { x } } =\frac { \frac { 1 }{ \cos ^{ 2 }{ x } } -\frac { \sin ^{ 2 }{ x } }{ \cos ^{ 2 }{ x } } }{ \frac { 1 }{ \sin { x } } } =\frac { \frac { 1-\sin ^{ 2 }{ x } }{ \cos ^{ 2 }{ x } } }{ \frac { 1 }{ \sin { x } } } =\frac { \frac { \cos ^{ 2 }{ x } }{ \cos ^{ 2 }{ x } } }{ \frac { 1 }{ \sin { x } } } =\frac { 1 }{ \frac { 1 }{ \sin { x } } } =\sin { x } \)

OpenStudy (imstuck):

Not such a big error! Good job!

OpenStudy (anonymous):

Thank you so much!!